Here are all the way to "waste" TP hits towards WSes, and exactly how much is wasted. Remember there is a WS
for 1 hit at the start, so >5 hits TPing is a waste.
3hit, 3hit - .2*.2 = 4% chance of occuring, 1 hit wasted
3hit, 1hit, 2hit - .2*.3*.5 = 3% chance of occuring, 1 hit wasted
3hit, 1hit, 3hit - .2*.3*.2 = 1.2% chance of occuring, 2hits wasted
2hit, 1hit, 3hit - .5*.3*.2 = 3% chance of occuring, 1 hit wasted
2hit, 2hit, 2hit - .5*.5*.5 = 12.5% chance of occuring, 1hit wasted
2hit, 2hit, 3hit - .5*.5*.2 = 5% chance of occuring, 2hits wasted
1hit, 2hit, 3hit - .3*.5*.2 = 3% chance of occuring, 1 hit wasted
1hit, 2hit, 1hit, 2hit - .3*.5*.3*.5 = 2.25% chance of occuring, 1hit wasted
1hit, 2hit, 1hit, 3hit - .3*.5*.3*.2 = .9% chance of occuring, 2hits wated
1hit, 3hit, 2hit - .3*.2*.5 = 3% chance of occuring, 1 hit wasted
1hit, 3hit, 3hit - .3*.2*.2 = 1.2% chance of occuring, 2hits wasted
1hit, 1hit, 2hit, 2hit - .3*.3*.5*.5 = 2.25% chance of occuring, 1hit wasted
1hit, 1hit, 2hit, 3hit - .3*.3*.5*.2 = .9% chance of occuring, 2hits wasted
1hit, 1hit, 1hit, 3hit - .3*.3*.3*.2 = .5% chance of occuring, 1hit wasted
1hit, 1hit, 1hit, 1hit, 2hit - .3*.3*.3*.3*.5 = .4% chance of occuring, 1hit wasted
1hit, 1hit, 1hit, 1hit, 3hit - .3*.3*.3*.3*.2 = .15% chance of occuring, 2hits wasted
Sum:
Chance to waste hits per TP phase:
4+3+3+12.5+3+2.25+3+2.25+.5+.4 = 33.9% chance to waste 1 hit
1.2+5+.9+1.2+.9+.15 = 9.35% chance to waste 2 hits
Average hits wasted per TP phase = .339*1+.0935*2 = .526
So, let's look at the WS ratios. Soboro swings 1.9 times as often as a Hagun, we'll say. So, when a Hagun
swings 5 times, a Soboro swings 9.5 times. Subtracting off .526 swings per 5, that becomes 9.5-.526*9.5/5 = 8.5
8.5 / 5 = 1.7 times the WSes as Hagun.